# FlowBotHD: History-Aware Diffuser Handling Ambiguities in Articulated Objects Manipulation

### v1

@@ Abstract
We introduce a novel approach to manipulate articulated objects with ambiguities, such as opening a door, in which multi-modality and occlusions create ambiguities about the opening side and direction. Multi-modality occurs when the method to open a fully closed door (push, pull, slide) is uncertain, or the side from which it should be opened is uncertain. Occlusions further obscure the door’s shape from certain angles, creating further ambiguities during the occlusion. To tackle these challenges, we propose a history-aware diffusion network that models the multi-modal distribution of the articulated object and uses history to disambiguate actions and make stable predictions under occlusions. Experiments and analysis demonstrate the state-of-art performance of our method and specifically improvements in ambiguity-caused failure modes. Our anonymized website is available at https://flowbothd.github.io/.
@@ 1 Introduction
Ambiguities are common in manipulation tasks. Consider opening a door as an example: when the door is fully closed, how it opens is ambiguous. Without further information, even experienced human beings can’t always tell by sight whether it opens by push, pull or even slide. Another type of ambiguity is caused by occlusions. When the door opens to certain angles, it might look like a line instead of a plane to a single view observer (see Figure 2). Occlusions cause loss of information, introducing ambiguities in the understanding of the object shape.
Faced with multi-modal ambiguities, a human might interact with the object and observe the outcome; for example, a person might push the door and see if it opens, and if not, they can adjust their strategy. Furthermore, momentary occlusions are not typically a problem for humans that can remember their past observations to supplement their current perception. Building on these insights, we expect an intelligent robot to have the ability to learn from multi-modal training data, predict multi-modal solutions for ambiguous observations, and with the help of historical observations, be able to make stable predictions even during occlusions.
To be generalizable and robust across different articulated object kinematics and geometries, we build on prior work that models articulated object manipulation with 3D articulation flow (3DAF) vectors as proposed in Eisner et al. [1]. This approach predicts a set of 3D vectors for each point on the articulated object that indicates its opening direction. Results from this approach have shown impressive abilities to generalize to novel objects in the real world, after training only in simulation. However, this approach fails when facing the ambiguous settings described above.
Therefore, in this paper, we propose a history-aware diffusion method to model multi-modality and handle occlusions. We use diffusion to model the multi-modal distribution and predict multiple possible object opening directions, and we use historical observations to adjust or disambiguate the distribution according to past trials. Our experiments and analysis demonstrate the benefits of our method in such ambiguous situations.
The contributions of this paper include:
@@
    A novel approach to handling both multi-modality ambiguity and occlusion ambiguity for articulated object manipulation.
@@
    A novel structure for incorporating object history.
@@ 2 Related Work
Articulated Object Manipulation: Manipulating articulated objects is an important and challenging task. One line of work investigated building manipulation systems with analytical methods [2, 3, 4, 5, 6]. With the help of large-scale datasets [7, 8], recent lines of work turn to learning-based methods. Several works model manipulation choices with explicit articulation parameters and poses [9, 10, 11, 12]. To improve the model’s generalization ability to different kinematics, several recent works learn to predict geometry-aware kinematics-free representations such as visual affordances [7, 13, 14] and articulation flows [1, 15] which achieve significant improvements in object generalization. Different from previous learning-based works which tend to use regression models for prediction, we propose to use diffusion models for better multi-modality handling.
@@
Uncertainty Handling for Robotics: Modeling uncertainty is a well-explored topic in various robotics fields [16, 17]. Previous works have included uncertainty in tasks including visual perception and tracking [18, 19, 20, 21], motion planning [22, 23, 24, 25], safety control [26] and also articulation parameter estimation [27] to build robust models. With similar motivation, our work focuses on building a model robust to visual uncertainty caused by multi-modality and occlusions.
@@
Diffusion Models in Robotics: Diffusion models, a class of generative models, learns to denoise an input to approach a given distribution. Following its impressive success in image generation, diffusion has also shown great potential in robotics tasks. Recent works have applied diffusion models to various tasks of robotics, including reinforcement learning [28, 29], imitation learning [30, 31, 32, 33, 34] and motion planning [35, 36, 37]. Similar to Ze et al. [34], the input of our diffusion model is a compact 3D point cloud which requires the model to have 3D awareness. Different from prior methods, we show how 3D diffusion models can be used for articulated object manipulation.
3 Background
3D Articulation Flow (3DAF):
@@ 3 Background
Suppose that an articulated object is described by a point cloud $P=\{p_{0},\ldots,p_{N-1}\}$ for each 3D point $p_{i}\in\mathbb{R}^{3}$ on the object visible surface. We assume that each point $p_{i}$ belongs to some part on the object, and that the parts are connected in a kinematic tree. We further assume that there are no closed-loop kinematic chains on the object that we are manipulating and also that each joint has only one degree of freedom. Following Eisner et al. [1], for each point $p_{i}$, we define the 3D Articulation Flow (3DAF) $f_{i}\in\mathbb{R}^{3}$ as the motion that point would undergo if there were an infinitesimal positive displacement $\delta\theta$ of the joint between that part and its parent. Eisner et al. [1] showed that, given a suction gripper, we can most efficiently open the articulated object using the predicted 3DAF by attaching the gripper to the point with the largest predicted flow norm $\arg\max_{i}\|f_{i}\|$ and then moving the gripper along its predicted flow direction $f_{i}$. The set of 3D Articulation Flows $F=\{f_{0},\ldots,f_{N-1}\}$ for all points in the point cloud is hierarchy-free (unlike a kinematic tree), and invariant under object translation and scaling, which allows for effective learning over a large set of objects of different categories and effective transfer from simulation to the real world.
Denoising Diffusion Probabilistic Models (DDPMs):
DDPMs [38] are a class of generative models that learn to progressively denoise towards the target data distribution starting from pure Gaussian noise. During the training process, for each ground-truth example, we iteratively add noise over multiple noising steps; we can compute the noised input in closed form as $x_{t}\sim\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ where $\bar{\alpha}_{t}$ is a noise parameter. The task objective is to train a noise predictor $\epsilon_{\theta}(x_{t},t)$ that inverts the added noise. We minimize the learning objective formulated as: $\mathbb{E}[{||\epsilon_{t}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},t)||_{2}}^{2}]$. With its iterative noise addition and denoising process, diffusion has demonstrated the ability to model complicated and multi-modal distributions.
4 Problem Statement and Assumptions
The task we propose to tackle is opening fully-closed objects from a single-view depth-sensor. The inputs to our model are a history of 3D point cloud observations of the articulated object being manipulated. As mentioned above, we assume that the articulated objects do not consist of a closed-loop kinematic chain and that each joint has only a single degree of freedom.
@@ 5 Method
As stated above, our goal is to open fully-closed articulated objects. To achieve this, we aim to predict how each articulated object might open, and we will use this prediction to determine how to manipulate the object. In making this prediction, we consider two types of ambiguity: multi-modality in how the object might open, and occlusion ambiguity, in which there is ambiguity from certain viewpoints because of object self-occlusions. The ambiguities will cause typical failure modes in existing methods[1]: the multi-modal ambiguity causes deterministic models to fail to open an object from a fully closed state; the occlusion ambiguity obscures models from making reasonable predictions from certain viewpoints, causing the robot to get stuck when the object passes through those viewpoints (see Figure 2). We aim to propose a robust model able to overcome the two failure modes and open articulated objects regardless of these ambiguities. To achieve this, we propose a history-aware diffusion model that handles ambiguity caused by both multi-modality and occlusion, together with a system for manipulating articulated objects based on our model’s predictions.
@@ 5.1 Diffusion for Multi-Modality
Due to the natural ambiguity of closed objects, there might be instances of objects with similar appearance but entirely different opening directions, such as a closed door that might open via push or pull. These examples cause multi-modality in both training datasets and in evaluation settings. To resolve these ambiguities, we take advantage of the ability of diffusion models to capture multi-modal distributions.
To condition the diffusion on a 3D point cloud input $\textbf{P}_{obs}$, we propose a novel 3D-aware diffusion structure. Our 3D-aware denoiser consists of a PointNet++ [39] encoder $\textbf{E}_{cur}$ that encodes the current inputs and a DiffusionTransformer (DiT) [40] denoiser D. During training, we first add noise to the ground truth flow $F^{*}$ (see Section 3) to obtain a noisy flow $F_{\epsilon}$, following the DDPM [38] diffusion formulation described in Section 3. Then we pair each point and its corresponding noisy flow as the current state $\textbf{S}_{cur}=\{\textbf{P}_{obs,i},F_{\epsilon,i}\},i\in[0,\textbf{M})$, where M is the total point count. We input the current state $\textbf{S}_{cur}$ through the point cloud encoder to obtain 3D geometry-aware point-wise features: $\textbf{E}_{cur}(\textbf{S}_{cur})$. Following the standard diffusion training method [38], we train the denoiser D conditioned on the current state encoding with the following learning objective:
$\mathbb{E}[{||\epsilon_{t}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},t,\textbf{E}_{cur}(\textbf{S}_{cur}))||_{2}}^{2}].$
@@ 5.2 Aggregating History for Disambiguation and Occlusion Handling
While the diffusion training method alone is capable of modeling multi-modal distributions, it lacks the ability to disambiguate based on past interactions. For example, during occlusions, if the robot succeeded in the past timesteps to partially open the door with the pull action, it should output a distribution that puts more probability on continuing to pull the door open in the future.
@@
Figure 1: FlowBotHD structure: the history observation and the history flow are input into a history encoder to obtain a global history latent. This history latent is injected into the current observation encoder through fully-connected projection layers and element-wise multiplication. The current observation encoder outputs a history-aware point-wise embedding. The history-aware embedding is then input to a DiT-based denoiser to predict the flow [1], i.e. the predicted future motion of the object if it were to be opened a small amount.
@@
Therefore, we propose to integrate the past history with a history encoder. We define a compact history state representation as the previous time step’s point cloud observation $\textbf{P}_{his}$, along with the previous step’s flow $F_{his}$. During training, we generate a synthetic history by sampling a random small angle to perturb the object to generate the “history” point cloud observation, and we obtain the corresponding ground truth articulation flow as the history flow. At inference time, we use the model’s previous point cloud observations and its previous flow predictions as history.
To aggregate history into the 3D-aware diffusion pipeline, we propose to fuse an embedding of the history state into the current state encoding process. Specifically, we first pair the history point cloud $\textbf{P}_{his}$ and the history flow $F_{his}$ into the history state $\textbf{S}_{his}=\{\textbf{P}_{his,i},f_{his,i}\},i\in[0,\textbf{M})$, where M is the total point count of the history step. Then we input $\textbf{S}_{his}$ to a PointNet++ [39] based history state encoder $\textbf{E}_{his}$ to obtain a global history latent embedding for the history point cloud: $\textbf{L}_{his}=\textbf{E}_{his}(\textbf{S}_{his})\in\mathbb{R}^{d_{h}}$, where $d_{h}$ is the dimension of the history latent.
To inject the history information into the current state encoder, we use a similar two-stream architecture to that of Shridhar et al. [41]; see Figure 1. In their case, the two streams are based on “what” and “where” whereas our two streams encode the current observation and the history. Consider a subset of layers in the encoder $\textbf{E}_{cur}$ from layer $k_{1}$ to layer $k_{2}$ as: {$e_{cur,k_{1}},\ldots,e_{cur,k_{2}}\}$, each of dimensionality $d_{k}$. We wish to element-wise multiply each of these layers by the history encoding $\textbf{L}_{his}$. In order to make the dimensionality match up, we define a series of linear projection layers $\textbf{W}_{k}\in\mathbb{R}^{d_{k}\cdot d_{h}}$. We then element-wise multiply (Hadmard product) each of the encoder layers with the projected history encoding: $e_{cur,k}^{\prime}=e_{cur,k}\odot(\textbf{W}_{k}\cdot\textbf{L}_{his}).$
To be compatible with situations at inference time where we don’t always have history available (such as the first timestep), we also include samples without providing history during training. To handle the absence of history, we use a learnable feature token $\textbf{L}_{0}$ as the history latent embedding for history injection. In summary, the history-aware encoder module $\textbf{E}_{ha}(\textbf{S}_{cur},\textbf{L}_{his})$ produces a history-aware encoding combining the current state observation and the history latent embedding:
$\textbf{F}_{ha}=\begin{cases}\textbf{E}_{ha}(\textbf{S}_{cur},\textbf{L}_{his}=\textbf{E}_{his}(\textbf{S}_{his}))&\text{if there is history},\\ \textbf{E}_{ha}(\textbf{S}_{cur},\textbf{L}_{his}=\textbf{L}_{0})&\text{otherwise}.\end{cases}$
Then we input the history-aware feature into the DiT-based denoiser D to output the predicted noise with learning objective: $\mathbb{E}[{||\epsilon_{t}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},t,\textbf{F}_{ha})||_{2}}^{2}].$ From the predicted noise, we are able to estimate the predicted flow $\hat{f}$ following the equations in Section 3.

### v2

@@ Abstract
We introduce a novel approach for manipulating articulated objects which are visually ambiguous, such doors which are symmetric or which are heavily occluded. These ambiguities can cause uncertainty over different possible articulation modes: for instance, when the articulation direction (e.g. push, pull, slide) or location (e.g. left side, right side) of a fully closed door are uncertain, or when distinguishing features like the plane of the door are occluded due to the viewing angle. To tackle these challenges, we propose a history-aware diffusion network that can model multi-modal distributions over articulation modes for articulated objects; our method further uses observation history to distinguish between modes and make stable predictions under occlusions. Experiments and analysis demonstrate that our method achieves state-of-art performance on articulated object manipulation and dramatically improves performance for articulated objects containing visual ambiguities. Our project website is available at https://flowbothd.github.io/.
@@
Figure 1: Our method, FlowBotHD, consists of a history-aware diffuser that handles multi-modality and occlusions in articulated object manipulation. Simulation and real world experiments show our model’s capability of opening general objects (including handling ambiguities and occlusions) as well as an ambiguous door.
@@ 1 Introduction
Ambiguities are common in manipulation tasks. Consider opening a door without a visible handle or hinges: when the door is fully closed, the way in which it opens may be totally ambiguous. It may swing open from the left side or the right side, and may swing either inwards or outwards. Without further information, even experienced humans can’t always tell by sight whether it opens by push, pull or even slide. Another type of ambiguity is caused by occlusions. When a door opens to certain angles such that the door’s principal plane is parallel to the viewing angle, it might look like a vertical line instead of a flat panel (see Figure 3). Occlusions cause loss of information, introducing ambiguities in the understanding of the object shape.
Faced with object ambiguities, a human might interact with the object and observe the outcome to disambiguate the true mode; for example, a person might push the door and see if it opens, and if not, they might try pulling the door open. Furthermore, momentary occlusions are not typically a problem for humans: we can remember past observations to supplement our current perception. We therefore would expect an intelligent robot to have the ability to learn from multi-modal training data, predict multi-modal solutions for ambiguous observations, and with the help of historical observations, make stable predictions even during occlusions.
To generalize across different articulated object kinematics and geometries, we build on prior work that models articulated object manipulation with 3D articulation flow (3DAF) vectors as proposed in Eisner et al. [1]. This approach predicts a set of 3D vectors for each point on the articulated object that indicates its opening direction. Results from this approach have shown impressive abilities to generalize to novel objects in the real world, after training only in simulation. However, since this prediction is unimodal and deterministic, this approach often fails to predict the correct flow when facing the ambiguous settings described above.
Therefore, in this paper, we propose a history-aware diffusion method to model multi-modality and handle occlusions for articulated objects. We use diffusion to model multi-modal distributions over multiple possible object opening directions, and we use historical observations to adjust or disambiguate the predicted distribution accordingly. Our experiments and analysis demonstrate the benefits of our method in such ambiguous situations. The contributions of this paper include:
@@
    A novel approach to model multi-modal ambiguity of articulated objects caused by visual ambiguity or occlusion.
@@
    A novel structure for incorporating object history to disambiguate articulation modes.
@@ 2 Related Work
Articulated Object Manipulation: Manipulating articulated objects is an important and challenging task. One line of work investigates building manipulation systems with analytical methods [2, 3, 4, 5, 6]. With the help of large-scale datasets [7, 8], recent lines of work turn to learning-based methods. Several works model manipulation choices with explicit articulation parameters and poses [9, 10, 11, 12]. To improve the model’s generalization ability to different kinematics, several recent works learn to predict geometry-aware kinematics-free representations such as visual affordances [7, 13, 14] and articulation flows [1, 15] which achieve significant improvements in object generalization. Different from previous learning-based works which tend to use regression models for prediction, we propose to use diffusion models for better handling of multi-modality.
@@
Uncertainty Modeling in Robotics: Modeling uncertainty is a well-explored topic in various robotics fields [16, 17]. Previous works have included uncertainty in tasks including visual perception and tracking [18, 19, 20, 21], motion planning [22, 23, 24, 25], safety control [26] and also articulation parameter estimation [27]. With similar motivation, our work focuses on building a model robust to visual uncertainty caused by multi-modality and occlusions.
@@
Diffusion Models in Robotics: Diffusion models, a class of generative models, learns to denoise an input to approach a given distribution. Following its impressive success in image generation, diffusion has also shown great potential in robotics tasks. Recent works have applied diffusion models to various tasks of robotics, including reinforcement learning [28, 29], imitation learning [30, 31, 32, 33, 34] and motion planning [35, 36, 37]. In contrast to prior methods, we show how 3D diffusion models can be used to handle ambiguities in articulated object manipulation.
3 Background: 3D Articulation Flow (3DAF)
@@
Our method builds upon the 3D Articulation Flow (3DAF) representation of Eisner et al. [1], which we briefly review here. We assume that the parts of the articulated object are connected in a kinematic tree with no closed-loop kinematic chains, and also that each joint has only one degree of freedom. Suppose that an articulated object is described by a point cloud $\mathbf{P}=\{p_{0},\ldots,p_{N-1}\}$ for a set of 3D points $p_{i}\in\mathbb{R}^{3}$ on the object visible surface. Following Eisner et al. [1], for each point $p_{i}$, we define the 3D Articulation Flow (3DAF) $f_{i}\in\mathbb{R}^{3}$ as the motion that point would undergo if there were an infinitesimal positive displacement $\delta\theta$ of the joint between that point’s articulation part and its parent. Eisner et al. [1] showed that, given a suction gripper, we can most efficiently open the articulated object using the predicted 3DAF by attaching the gripper to the point $p_{i_{\text{max}}}$ with the largest predicted flow norm $i_{\text{max}}=\arg\max_{i}\|f_{i}\|$ and then moving the gripper along its predicted flow direction. The set of 3D Articulation Flows $\mathbf{F}=\{f_{0},\ldots,f_{N-1}\}$ for all points in the point cloud is hierarchy-free and invariant under object translation and scaling, which allows for effective learning over a large set of objects of different categories and effective transfer from simulation to the real world.
@@ 4 Method
The task we propose to tackle is opening fully-closed articulated objects from a single-view depth-sensor. The inputs to our model are a history of 3D point cloud observations of the articulated object being manipulated. We assume that the articulated objects do not consist of a closed-loop kinematic chain and that each joint has only a single degree of freedom.
We aim to enable a robot to manipulate ambiguous articulated objects, focusing on two types of ambiguity: ambiguity in opening direction due to lack of visual cues, and ambiguity due to self-occlusion during articulation. Both types of ambiguity are known failure modes of prior deterministic systems [1]. At a high level, our objective is to predict all the modes that describe how a potentially-ambiguous articulated object might open, and then to use these predictions to determine how to manipulate the object. Below we will describe our proposed history-aware diffusion model which can represent multimodality (Section 4.1) and is robust to occlusion (Section 4.2), as well as a system for manipulating articulated objects based on multimodal predictions (Section 4.3).
@@ 4.1 Diffusion for Multi-Modality
Our system builds on the 3D Articulation Flow representation from prior work [1] described in Section 3. Our first task is to construct a system which can accurately predict multi-modal affordances for an articulated object, in which affordances are captured by 3D Articulation Flow. For instance, given an ambiguous door that can open right or left (as shown in Figure 4), we’d like to be able to generate one set of flow vectors to describe the possibility that the door opens rightward, and a different set of flow vectors to capture the possibility that it might open leftward. Diffusion processes have been shown to be a good choice of model architecture for this setting, as they can represent multi-modal distributions over high-dimensional data [38, 30].
We now describe how we can use diffusion to predict multi-modal 3D Articulation flow. To incorporate observed 3D point cloud inputs $\mathbf{P}_{\text{obs}}$ into the standard diffusion model paradigm, we use a PointNet++ [39] encoder $g_{\phi}$ that encodes the current inputs and a DiffusionTransformer (DiT) [40] denoiser $\mathbf{D}_{\theta}$. During training, we first add noise to the ground truth flow $\mathbf{F}_{\text{obs}}^{*}$ to obtain a noisy flow $\mathbf{F}_{\text{obs}}^{\epsilon}$, following the DDPM [38] diffusion formulation. Then we pair each point and its corresponding noisy flow as the current state $\mathbf{S}_{\text{cur}}=\{p_{\text{obs},i},f^{\epsilon}_{\text{obs},i}\},i\in[1\dots N]$, where $N$ is the total number of observed points in $\mathbf{P}_{\text{obs}}$. We encode the current state $\mathbf{S}_{\text{cur}}$ as $\Psi_{\text{obs}}=g_{\phi}(\mathbf{S}_{\text{cur}})$ to obtain 3D geometry-aware point-wise features. Following the standard diffusion training procedure [38], we train the denoiser $\mathbf{D}_{\theta}$ and encoder $g_{\phi}$ with the following learning objective:
$\mathcal{J}(\theta,\phi)=\mathbb{E}[{||\epsilon_{t}-\mathbf{D}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},t,g_{\phi}(\mathbf{S}_{\text{cur}}))||_{2}^{2}}]$
From the predicted noise, we estimate the flow $\hat{\mathbf{F}}$ following the standard diffusion equations [38].
@@ 4.2 Aggregating History for Disambiguation and Occlusion Handling
While this proposed 3DAF diffusion model is capable of modeling multi-modal distributions, it lacks the ability to disambiguate based on past interactions. For example, suppose the door is occluded in a current observation, as shown in Figure 3; if the robot succeeded in the past timesteps to partially open the door, it should output a distribution that places more probability on a flow which agrees with the previously observed motion.
@@
Figure 2: FlowBotHD structure: a history observation and the history flow are input into a history encoder to obtain a global history latent. This history latent is injected into the current observation encoder through fully-connected projection layers and element-wise multiplication. The current observation encoder outputs a history-aware point-wise embedding. The history-aware embedding is then input to a DiT-based denoiser to predict the 3D articulation flow [1], i.e. the predicted future motion of the object if it were to be opened a small amount.
To accomplish this objective, we propose to modify the above architecture to integrate the past history with a history encoder. We define a compact history state representation as the previous time step’s point cloud observation $\textbf{P}_{\text{his}}$, along with the previous step’s flow $\mathbf{F}_{\text{his}}$, as $\mathbf{S}_{\text{his}}=\{p_{\text{his},i},f_{\text{his},i}\},i\in[1\dots M]$, where $M$ is the number of points in the previous point cloud. We then pass $\mathbf{S}_{\text{his}}$ though a global point cloud encoder $h_{\gamma}$ based on PointNet++ [39] to get a global history context vector $z_{\text{his}}=h_{\gamma}(\mathbf{S}_{\text{his}})\in\mathbb{R}^{d_{\text{his}}}$, where $d_{\text{his}}$ is the fixed dimension of the history latent. We then inject this additional context into the “decoding” layers of our U-net encoder $g_{\phi}$, similar to the method proposed by Shridhar et al. [41] (see Figure 2). For each decoding layer $j$ of $g_{\phi}$ with point-wise activations $a_{j}\in\mathbb{R}^{d_{j}}$, we apply a learned linear projection $\mathbf{W}_{j}\in\mathbb{R}^{d_{\text{his}}\times d_{j}}$ to the history latent $z_{\text{his}}$, and combine with the original activation using the Hadamard product: $a_{j}^{\prime}=a_{j}\odot(\mathbf{W}_{j}\cdot z_{\text{his}})$.
@@
During training, we generate a synthetic history by sampling a random small angle to perturb the object to generate the “history” point cloud observation, and we obtain the corresponding ground truth articulation flow as the history flow. At inference time, we use the model’s previous point cloud observations and its previous flow predictions as history. To account for situations at inference time when no history yet exists (i.e. the first timestep), we also train with samples that include no history, and include a learnable “start” vector $z_{0}$ in place of the history vector $z_{\text{his}}$ in such cases. We train this model end-to-end with the standard diffusion objective outlined in Section 4.1.